OpenStudy (anonymous):

Solve the equation x ^1/4 = -4x^1/2 + 22. Plot both sides of the equation in the same viewing screen, y1= x^1/4 and y2=-4x^1/2+22. Does the point(s) of intersection agree with your solution?

OpenStudy (anonymous):

let u=x^(1/4)

OpenStudy (anonymous):

and u^2=x^(1/2)

OpenStudy (anonymous):

then solve it as quadratic equation.

OpenStudy (anonymous):

u=-4u^2+22

OpenStudy (anonymous):

are you with me?

OpenStudy (anonymous):

4u^2+u-22=0

OpenStudy (anonymous):

use Quadratic Formula to solve it http://prntscr.com/8am6wc

OpenStudy (anonymous):

this the final answer try to find it x= \[\frac{ ({\sqrt{353}-1})^4 }{ 4096}\]

OpenStudy (anonymous):

thanks sir ur the best